# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ V.2 Gauge issues, effective action, and stability conditions
$-\overline{J}\overline{\rho}_{m,n}\dot{\delta\mathcal{B}}_{p}+\left(\frac{\overline{J}}{a^{2}}\overline{\rho}_{m,n}+\frac{1-\xi b_{t}^{2}}{2\kappa a}k_{z}^{2}\right)\lambda_{p}-\frac{\xi b_{t}}{2\kappa a}k_{z}^{2}\varsigma_{p}=0,$ (71)
@@
$S^{(2)}_{v}=\int dtd^{3}x$ $\Bigg[$ $\frac{a}{2}\delta^{pq}\dot{\varsigma}_{p}\dot{\varsigma}_{q}-\left(\frac{1}{2a}+\frac{\xi^{2}b_{t}^{2}k_{z}^{2}}{4\kappa(1-\xi b_{t}^{2})ak_{z}^{2}+8\kappa^{2}\overline{J}\overline{\rho}_{m,n}}\right)k_{z}^{2}\delta^{pq}\varsigma_{p}\varsigma_{q}-\frac{\xi a}{\kappa}\dot{H}\delta^{pq}\varsigma_{p}\varsigma_{q}$ (72)
$+\frac{\overline{J}(1-\xi b_{t}^{2})a^{3}\overline{\rho}_{m,n}k_{z}^{2}}{2(1-\xi b_{t}^{2})ak_{z}^{2}+4\kappa\overline{J}\overline{\rho}_{m,n}}\delta^{pq}\dot{\delta\mathcal{B}}_{p}\dot{\delta\mathcal{B}}_{q}-\frac{\xi b_{t}\overline{J}a\overline{\rho}_{m,n}k_{z}^{2}}{(1-\xi b_{t}^{2})ak_{z}^{2}+2\kappa\overline{J}\overline{\rho}_{m,n}}\delta^{pq}\varsigma_{p}\dot{\delta\mathcal{B}}_{q}\;\;\Bigg].$
@@
$\frac{\overline{J}a\overline{\rho}_{m,n}k_{z}^{2}}{(1-\xi b_{t}^{2})ak_{z}^{2}+2\kappa\overline{J}\overline{\rho}_{m,n}}\big(\xi b_{t}^{2}\varsigma_{p}-(1-\xi b_{t}^{2})a^{2}\dot{\delta\mathcal{B}}_{p}\big)=C^{v}_{p},$ (73)
where $C^{v}_{p}$ are constants. Since we are only concerned with the dynamic effects, we can choose $C_{p}=0$ without altering the physics. For a more general case, we can solve for $C_{p}$, which depends non-locally on $\varsigma_{p}$ (see Ref. Dyer:2008hb ). However, we are not concerned with this general case. Substituting this equation (73) into the action (72) and eliminating $\dot{\delta\mathcal{B}}_{p}$, we obtain an effective action containing only dynamical variables
@@
where $c^{2}_{v}=1+\xi^{2}b_{t}^{2}/\big(2\kappa(1-\xi b_{t}^{2})\big)$ and $m^{2}_{v}=2\xi\dot{H}/\kappa$. $c^{2}_{v}$ and $m^{2}_{v}$ are the vector propagation speed squared and effective mass squared, respectively. It is easy to see that there are two dynamical degrees of freedom $\varsigma_{1}$ and $\varsigma_{2}$ for the vector perturbations. In Fourier space, this action is equivalent with the original one under the gauge condition (70) when the constraints (54), (57), (71), and (73) are considered (with $C_{p}=0$). When the boundary conditions are specified, these constraints imply that the non-dynamical variables $\lambda_{p}$, $\varepsilon_{p}$, $\chi^{p}$, $\delta\mathcal{A}_{p}$, and $\delta\mathcal{B}_{p}$ all depend on the variable $\varsigma_{p}$.
@@
The condition $|\xi b_{t}^{2}|\ll 1$ for the Lorentz-violating parameter ensures Laplacian stability. To avoid tachyonic instability, and considering that $\dot{H}<0$ (see Eq. (29)), the following requirement must be met
@@
Combining this condition with $\overline{V}_{,b^{2}}/\xi>0$ (see Eq. (30)), the free of tachyonic instability implies $\overline{V}_{,b^{2}}(t_{0})\leq 0$, which imposes a further constraint on the potential term of the action.
@@
where $w_{v}$ is the frequency of the vector perturbations. Since $\kappa>0$ and $|\xi b_{t}^{2}|\ll 1$, the propagation speed of the vector perturbations is greater than 11 $\big(c_{v}=1+(\xi/4\kappa)\xi b_{t}^{2}+\mathcal{O}\big((\xi b_{t}^{2})^{2}\big)\big)$, which is caused by the violation of Lorentz symmetry. If the Lorentz-violating parameter $\xi b_{t}^{2}$ equals zero, the speed will return to $1$.
@@ VI.1 The second-order action of the scalar perturbations
$S^{(2)}_{s}=\int dtd^{3}x$ $\Bigg[$ $-\frac{a^{3}}{2\kappa}\left(3(2-\xi b_{t}^{2})H^{2}+9\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\right)\phi_{h}^{2}+\frac{\xi b_{t}^{2}}{\kappa}a(\partial\phi_{h})^{2}-2b_{t}^{2}\overline{V}_{,b^{2}b^{2}}a^{3}\phi_{b}^{2}$ (84)
$+\frac{a}{2}(\partial\phi_{b})^{2}+\left(\frac{b_{t}}{\kappa}a^{3}(-3\xi H^{2}+3\xi\dot{H}+4\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}})\phi_{b}+\frac{\xi b_{t}}{\kappa}a\partial^{2}\phi_{b}-\frac{3\xi b_{t}}{\kappa}a^{3}H\dot{\phi}_{b}\right)\phi_{h}$
$-\frac{\overline{\rho}_{m,nn}}{2a^{3}}\phi_{m}^{2}+\frac{\overline{\rho}_{m,n}}{2\overline{J}a^{2}}(\partial\varphi_{m})^{2}-(\overline{\rho}_{m,n}\phi_{h}+\dot{\phi}_{\ell})\phi_{m}+\frac{1}{a^{2}}\varphi_{m}\partial^{2}\phi_{\ell}+\mathcal{L}_{\alpha}+\mathcal{L}_{E}+\mathcal{L}_{\varphi_{b}}+\mathcal{L}_{\varphi_{h}}\;\;\Bigg].$
@@
$\mathcal{L}_{\alpha}$ $=$ $-\frac{\overline{J}^{2}\overline{\rho}_{m,nn}}{8a^{3}}(\partial^{2}\alpha)^{2}+\frac{\xi b_{t}}{2\kappa}a^{3}(b_{t}\partial^{2}\phi_{h}-\partial^{2}\phi_{b})\ddot{\alpha}+\frac{a^{3}}{\kappa}H(\partial^{2}\phi_{h}-\xi b_{t}\partial^{2}\phi_{b})\dot{\alpha}$ (85)
$-\left(\frac{1-\xi b_{t}^{2}}{\kappa}a^{3}\dot{H}\partial^{2}\phi_{h}+\frac{3\overline{J}^{2}\overline{\rho}_{m,nn}}{4a^{3}}\partial^{2}E-\frac{1-\xi b_{t}^{2}}{2\kappa}\partial^{2}\partial_{t}(a^{3}\dot{E})-\frac{\overline{J}\overline{\rho}_{m,nn}}{2a^{3}}\partial^{2}\phi_{m}\right)\alpha,$
$\mathcal{L}_{E}$ $=$ $\frac{a}{4\kappa}(\partial E)^{2}-\frac{3(1-\xi b_{t}^{2})}{4\kappa}a^{3}\dot{E}^{2}-\frac{9\overline{J}^{2}\overline{\rho}_{m,nn}}{8a^{3}}E^{2}-\left(\frac{3(1-\xi b_{t}^{2})}{\kappa}a^{3}\dot{H}\phi_{h}+\frac{a}{\kappa}\partial^{2}\phi_{h}-\frac{3\overline{J}\overline{\rho}_{m,nn}}{2a^{3}}\phi_{m}\right)E$ (86)
$+\frac{1}{\kappa}\left(3a^{3}H(\phi_{h}-\xi b_{t}\phi_{b})+(1-\xi b_{t}^{2})a\partial^{2}\varphi_{h}-\xi b_{t}a\partial^{2}\varphi_{b}\right)\dot{E}+\frac{3\xi b_{t}}{2\kappa}a^{3}(b_{t}\phi_{h}-\phi_{b})\ddot{E},$
@@
$\mathcal{L}_{\varphi_{h}}$ $=$ $-\frac{1-\xi b_{t}^{2}}{\kappa}a\dot{H}(\partial\varphi_{h})^{2}-\frac{\xi b_{t}}{\kappa}a(b_{t}\partial^{2}\phi_{h}-\partial^{2}\phi_{b})\dot{\varphi}_{h}-\left(\frac{2(1-\xi b_{t}^{2})}{\kappa}aH\partial^{2}\phi_{h}+\frac{\overline{\rho}_{m,n}}{a^{2}}\partial^{2}\varphi_{m}\right)\varphi_{h}.$ (88)
@@
VI.2 Gauge I: $\alpha=0,\;E=0$
@@
Varying the action (84) with respect to $\varphi_{m}$, $\phi_{\ell}$, and $\phi_{h}$, respectively, we can obtain the constraints for them in Fourier space
@@
$\phi_{h}=\Big(\xi b_{t}^{2}k^{2}\dot{\varphi}_{h}+2(1-\xi b_{t}^{2})Hk^{2}\varphi_{h}-3\xi b_{t}a^{2}H\dot{\phi}_{b}-\xi b_{t}k^{2}\phi_{b}-b_{t}a^{2}(3\xi H^{2}-3\xi\dot{H}-4\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}})\phi_{b}$
$\qquad\quad-2\xi b_{t}Hk^{2}\varphi_{b}-\frac{\kappa}{a}\overline{\rho}_{m,n}\phi_{m}\Big)/(D_{1}-2\xi b_{t}^{2}k^{2}),$ (95)
@@
where $k$ is a wavenumber, $k^{2}=\delta^{ij}k_{i}k_{j}$, and $D_{1}=3(2-\xi b_{t}^{2})a^{2}H^{2}+9\xi b_{t}^{2}a^{2}\dot{H}+4\kappa b_{t}^{4}a^{2}\overline{V}_{,b^{2}b^{2}}$. By substituting these three constraints one by one into the action (84) in Fourier space, we can eliminate the variables $\varphi_{m}$, $\phi_{\ell}$, and $\phi_{h}$. In order to eliminate the coupling terms between the time derivatives of the variables, which greatly facilitates the identification of non-dynamic variables, we define a new variable
$\psi_{1}\equiv\varphi_{h}-\frac{3a^{2}}{b_{t}k^{2}}H\phi_{b}.$ (96)
@@
The variable $\varphi_{h}$ can then be eliminated by $\psi_{1}$, yielding an effective action. This action is equivalent to the original one. Since it serves merely as an auxiliary action, we will not present its specific form here. Likewise, any subsequent auxiliary actions will also be omitted.
@@
By analyzing the newly obtained action, one can determine whether any non-dynamical perturbations remain by checking if all perturbations possess kinetic terms. In our newly obtained action, the perturbation $\phi_{b}$ leads to a constraint
$\phi_{b}=\frac{1}{F_{1}(k,6)}\left(F_{2}(k,6)\dot{\varphi}_{b}+F_{3}(k,6)\varphi_{b}+F_{4}(k,6)\dot{\psi}_{1}+F_{5}(k,6)\psi_{1}+F_{6}(k,2)\dot{\phi}_{m}+F_{7}(k,4)\phi_{m}\right).$ (97)
Here, $F_{n}(k,m)$ are functions of $t$, the index $n$ labels distinct functions, and $m$ denotes the highest power of the wavenumber $k$ in $F_{n}(k,m)$. The specific forms of these coefficients are provided in the appendix A.
@@
Substituting the constraint (97) into the newly obtained action allows us to eliminate the variable $\phi_{b}$. Subsequently, we define two new variables to eliminate the coupling terms between the time derivatives of the remaining variables,
$\psi_{2}\equiv\psi_{1}+\left(\frac{6(1-\xi b_{t}^{2})a^{2}H^{2}}{\kappa b_{t}k^{2}}+\frac{b_{t}k^{2}}{k^{2}-3a^{2}\dot{H}}\right)^{-1}\varphi_{b}-\frac{3\kappa aH(k^{2}-3a^{2}\dot{H})\overline{\rho}_{m,n}}{\kappa b_{t}^{2}k^{6}+6(1-\xi b_{t}^{2})(k^{2}-3a^{2}\dot{H})a^{2}H^{2}k^{2}}\phi_{m},$ (98)
$\psi_{3}\equiv\varphi_{b}+\frac{\kappa b_{t}\overline{\rho}_{m,n}}{2(1-\xi b_{t}^{2})(k^{2}-3a^{2}\dot{H})aH}\phi_{m}.$ (99)
@@
Using these two new variables, we can eliminate $\psi_{1}$ and $\varphi_{b}$ from the action. The resulting action can be written as
$S^{(2)}_{s}=\int dtd^{3}x$ $\Bigg[$ $-\frac{\xi b_{t}^{2}a}{2(2\kappa-\xi)}(k\dot{\psi}_{2})^{2}+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}+\frac{3(1-\xi b_{t}^{2})}{\kappa b_{t}^{2}}a^{3}H^{2}(\dot{\psi}_{3})^{2}-\frac{\overline{\rho}_{m,nn}}{2a^{3}}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}$ (100)
$-\frac{2\kappa+\xi b_{t}^{2}(\xi-2\kappa)}{\kappa\xi b_{t}^{4}(2\kappa-\xi)}aH^{2}k^{2}\psi_{3}^{2}-\frac{2(\kappa-2\kappa\xi b_{t}^{2}+\xi^{2}b_{t}^{2})}{\kappa b_{t}(2\kappa-\xi)}aHk\psi_{3}(k\dot{\psi}_{2})+M_{1}\left(k\psi_{2},\frac{1}{k}\phi_{m},\psi_{3}\right)\;\;\Bigg],$
where $M_{1}$ encompasses all terms of order lower than $k^{2}$. In this action, it is $\dot{\psi}_{2}$, and not $\psi_{2}$ itself, that contributes to the $k^{2}$-order terms associated with the variable $\psi_{2}$.
In the small-scale limit ($k\rightarrow\infty$), varying the action (100) with respect to $\psi_{2}$ yields the following equation of motion (analogous to the treatment of the vector perturbations)
$k\dot{\psi}_{2}\approx-\frac{2}{\kappa\xi b_{t}^{3}}(\kappa-2\kappa\xi b_{t}^{2}+\xi^{2}b_{t}^{2})Hk\psi_{3}.$ (101)
Here, we consider only the dynamical effects and have accordingly simplified the expression. In the small-scale limit, substituting this constraint into the action (100), an effective action that contains only the dynamical variables is obtained
$S^{(2)}_{s}\approx\int dtd^{3}x\left[\frac{3(1-\xi b_{t}^{2})}{\kappa b_{t}^{2}}a^{3}H^{2}\left((\dot{\psi}_{3})^{2}-c_{s}^{2}\frac{k^{2}}{a^{2}}\psi_{3}^{2}\right)+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}\left(\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}-\frac{\overline{p}_{m,n}}{\overline{\rho}_{m,n}}\frac{k^{2}}{a^{2}}\left(\frac{\phi_{m}}{k}\right)^{2}\right)\right],$ (102)
where $c_{s}^{2}\equiv(3\kappa+2\xi b_{t}^{2}(\xi-2\kappa))/\big(3\kappa(1-\xi b_{t}^{2})\big)$. It is straightforward to see that the variables $\phi_{h}$, $\varphi_{h}$, $\phi_{b}$, $\varphi_{b}$, $\varphi_{m}$, and $\phi_{\ell}$ depend on $\psi_{2}$, $\psi_{3}$, and $\phi_{m}$, as shown in Eqs. (93)-(99). The action (100) indicates that $\psi_{2}$ is a cyclic coordinate in the small-scale limit, and the constraint (101) implies that $\dot{\psi}_{2}$ depends on $\psi_{3}$.

### v3

@@ V.2 Gauge issues, effective action, and stability conditions
$-\overline{J}\overline{\rho}_{m,n}\dot{\delta\mathcal{B}}_{p}+\left(\frac{\overline{J}}{a^{2}}\overline{\rho}_{m,n}+\frac{1-(\xi+\sigma)b_{t}^{2}}{2\kappa a}k_{z}^{2}\right)\lambda_{p}-\frac{\xi b_{t}}{2\kappa a}k_{z}^{2}\varsigma_{p}=0,$ (70)
@@
$S^{(2)}_{v}=\int dtd^{3}x$ $\Bigg[$ $\frac{a}{2}\delta^{pq}\dot{\varsigma}_{p}\dot{\varsigma}_{q}-\left(\frac{1}{2a}+\frac{\xi^{2}b_{t}^{2}k_{z}^{2}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)ak_{z}^{2}+8\kappa^{2}\overline{J}\overline{\rho}_{m,n}}\right)k_{z}^{2}\delta^{pq}\varsigma_{p}\varsigma_{q}-\frac{\xi a}{\kappa}\dot{H}\delta^{pq}\varsigma_{p}\varsigma_{q}$ (71)
$+\frac{\overline{J}\big(1-(\xi+\sigma)b_{t}^{2}\big)a^{3}\overline{\rho}_{m,n}k_{z}^{2}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)ak_{z}^{2}+4\kappa\overline{J}\overline{\rho}_{m,n}}\delta^{pq}\dot{\delta\mathcal{B}}_{p}\dot{\delta\mathcal{B}}_{q}-\frac{\xi b_{t}\overline{J}a\overline{\rho}_{m,n}k_{z}^{2}}{\big(1-(\xi+\sigma\big)b_{t}^{2})ak_{z}^{2}+2\kappa\overline{J}\overline{\rho}_{m,n}}\delta^{pq}\varsigma_{p}\dot{\delta\mathcal{B}}_{q}\;\;\Bigg].$
@@
$\frac{\overline{J}a\overline{\rho}_{m,n}k_{z}^{2}}{\big(1-(\xi+\sigma)b_{t}^{2}\big)ak_{z}^{2}+2\kappa\overline{J}\overline{\rho}_{m,n}}\left(\xi b_{t}\varsigma_{p}-\big(1-(\xi+\sigma)b_{t}^{2}\big)a^{2}\dot{\delta\mathcal{B}}_{p}\right)=C^{v}_{p},$ (72)
where $C^{v}_{p}$ are constants. Since we are only concerned with the dynamic effects, we can choose $C_{p}^{v}=0$ without altering the physics. For a more general case, we can solve for $C_{p}^{v}$, which depends non-locally on $\varsigma_{p}$ (see Ref. Dyer and Hinterbichler (2009)). However, we are not concerned with this general case. Substituting this equation (72) into the action (71) and eliminating $\dot{\delta\mathcal{B}}_{p}$, we obtain an effective action containing only dynamical variables
@@
where $c^{2}_{v}=1+\xi^{2}b_{t}^{2}/\big(2\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)\big)$ and $m^{2}_{v}=2\xi\dot{H}/\kappa$. $c^{2}_{v}$ and $m^{2}_{v}$ are the vector propagation speed squared and effective mass squared, respectively. It is easy to see that there are two dynamical degrees of freedom $\varsigma_{1}$ and $\varsigma_{2}$ for the vector perturbations. In Fourier space, this action is equivalent with the original one under the gauge condition (69) when the constraints (53), (56), (70), and (72) are considered (with $C_{p}^{v}=0$). When the boundary conditions are specified, these constraints imply that the non-dynamical variables $\lambda_{p}$, $\varepsilon_{p}$, $\chi^{p}$, $\delta\mathcal{A}_{p}$, and $\delta\mathcal{B}_{p}$ all depend on the variable $\varsigma_{p}$.
@@
The condition $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$ for the Lorentz-violating parameters ensure Laplacian stability. To avoid tachyonic instability, and considering that $\dot{H}<0$ (see Eq. (29)), the following requirement must be met
@@
This condition further tightens the bound on the speed of the tensor GW in the Bumblebee model, see Eq. (48).
@@
where $w_{v}$ is the frequency of the vector perturbations. Since $\kappa>0$ and $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, the propagation speed of the vector perturbations is greater than $1$, which is caused by the violation of Lorentz symmetry. If the Lorentz-violating parameter $\xi b_{t}^{2}$ equals zero, the speed will return to $1$.
@@ VI.1 The second-order action of the scalar perturbations
$S^{(2)}_{s}=\int dtd^{3}x$ $\Bigg[$ $-\frac{a^{3}}{2\kappa}\left(3\big(2-(\xi-8\sigma)b_{t}^{2}\big)H^{2}+9(\xi+2\sigma)b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\right)\phi_{h}^{2}+\frac{(\xi+2\sigma)b_{t}^{2}}{\kappa}a(\partial\phi_{h})^{2}$ (83)
$+\left(\frac{b_{t}}{\kappa}a^{3}\big(-3(\xi-2\sigma)H^{2}+3(\xi+2\sigma)\dot{H}+4\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}}\big)\phi_{b}+\frac{(\xi+2\sigma)b_{t}}{\kappa}a\partial^{2}\phi_{b}-\frac{3(\xi+2\sigma)b_{t}}{\kappa}a^{3}H\dot{\phi}_{b}\right)\phi_{h}$
$-2b_{t}^{2}\overline{V}_{,b^{2}b^{2}}a^{3}\phi_{b}^{2}+\frac{a}{2}(\partial\phi_{b})^{2}-\frac{\overline{\rho}_{m,nn}}{2a^{3}}\phi_{m}^{2}+\frac{\overline{\rho}_{m,n}}{2\overline{J}a^{2}}(\partial\varphi_{m})^{2}-(\overline{\rho}_{m,n}\phi_{h}+\dot{\phi}_{\ell})\phi_{m}+\frac{1}{a^{2}}\varphi_{m}\partial^{2}\phi_{\ell}$
$+\mathcal{L}_{\alpha}+\mathcal{L}_{E}+\mathcal{L}_{\varphi_{b}}+\mathcal{L}_{\varphi_{h}}\;\;\Bigg].$
@@
$\mathcal{L}_{\alpha}$ $=$ $-\frac{\overline{J}^{2}\overline{\rho}_{m,nn}}{8a^{3}}(\partial^{2}\alpha)^{2}+\frac{(\xi+2\sigma)b_{t}}{2\kappa}a^{3}\big(b_{t}\partial^{2}\phi_{h}-\partial^{2}\phi_{b}\big)\ddot{\alpha}+\frac{a^{3}}{\kappa}H\big((1+3\sigma b_{t}^{2})\partial^{2}\phi_{h}-(\xi+4\sigma)b_{t}\partial^{2}\phi_{b}\big)\dot{\alpha}$ (84)
$-\left(\frac{1-(\xi+\sigma)b_{t}^{2}}{\kappa}a^{3}\dot{H}\partial^{2}\phi_{h}+\frac{3\overline{J}^{2}\overline{\rho}_{m,nn}}{4a^{3}}\partial^{2}E-\frac{1-(\xi+\sigma)b_{t}^{2}}{2\kappa}\partial^{2}\partial_{t}(a^{3}\dot{E})-\frac{\overline{J}\overline{\rho}_{m,nn}}{2a^{3}}\partial^{2}\phi_{m}\right)\alpha,$
$\mathcal{L}_{E}$ $=$ $\frac{(1-\sigma b_{t}^{2})a}{4\kappa}(\partial E)^{2}-\frac{3\big(1-(\xi+\sigma)b_{t}^{2}\big)}{4\kappa}a^{3}\dot{E}^{2}-\frac{9\overline{J}^{2}\overline{\rho}_{m,nn}}{8a^{3}}E^{2}+\frac{3(\xi+2\sigma)b_{t}}{2\kappa}a^{3}(b_{t}\phi_{h}-\phi_{b})\ddot{E}$ (85)
$-\left(\frac{3\big(1-(\xi+\sigma)b_{t}^{2}\big)}{\kappa}a^{3}\dot{H}\phi_{h}+\frac{(1+\sigma b_{t}^{2})a}{\kappa}\partial^{2}\phi_{h}-\frac{2\sigma b_{t}a}{\kappa}\partial^{2}\phi_{b}-\frac{3\overline{J}\overline{\rho}_{m,nn}}{2a^{3}}\phi_{m}\right)E$
$+\frac{1}{\kappa}\left(3a^{3}H\big((1+3\sigma b_{t}^{2})\phi_{h}-(\xi+4\sigma)b_{t}\phi_{b}\big)+\big(1-(\xi+\sigma)b_{t}^{2}\big)a\partial^{2}\varphi_{h}-\xi b_{t}a\partial^{2}\varphi_{b}\right)\dot{E},$
@@
$\mathcal{L}_{\varphi_{h}}$ $=$ $-\frac{1-(\xi+\sigma)b_{t}^{2}}{\kappa}a\dot{H}(\partial\varphi_{h})^{2}-\left(\frac{2(1-(\xi-\sigma)b_{t}^{2})}{\kappa}aH\partial^{2}\phi_{h}-\frac{4\sigma b_{t}}{\kappa}aH\partial^{2}\phi_{b}+\frac{\overline{\rho}_{m,n}}{a^{2}}\partial^{2}\varphi_{m}\right)\varphi_{h}$ (87)
$-\frac{(\xi+2\sigma)b_{t}}{\kappa}a(b_{t}\partial^{2}\phi_{h}-\partial^{2}\phi_{b})\dot{\varphi}_{h}.$
@@
VI.2 Effective action and stability conditions
@@
Varying the action (83) with respect to $\varphi_{m}$, $\phi_{\ell}$, and $\varphi_{h}$, respectively, we can obtain the constraints for them in Fourier space
@@
$\varphi_{h}=\Big((2-3\xi b_{t}^{2})H\phi_{h}-(\xi+2\sigma)b_{t}^{2}\dot{\phi}_{h}+(\xi-2\sigma)b_{t}H\phi_{b}+(\xi+2\sigma)b_{t}\dot{\phi}_{b}+\frac{\kappa\overline{\rho}_{m,n}}{a^{3}}\varphi_{m}\Big)/\Big(2\dot{H}\big(1-(\xi+\sigma)b_{t}^{2}\big)\Big),$ (94)
@@
where $k$ is a wavenumber, $k^{2}=\delta^{ij}k_{i}k_{j}$. By substituting these three constraints one by one into the action (83) in Fourier space, we can eliminate the variables $\varphi_{m}$, $\phi_{\ell}$, and $\varphi_{h}$. In order to eliminate the coupling terms between the time derivatives of the variables, which greatly facilitates the identification of non-dynamic variables, we define a new variable
$\psi_{1}\equiv\phi_{b}-b_{t}\phi_{h}+\frac{\kappa\overline{\rho}_{m,n}}{(\xi+2\sigma)b_{t}ak^{2}}\phi_{m}.$ (95)
@@
The variable $\phi_{b}$ can then be eliminated by $\psi_{1}$, yielding an effective action. This action is equivalent to the original one. Since it serves merely as an auxiliary action, we will not present its specific form here. Likewise, any subsequent auxiliary actions will also be omitted.
@@
By analyzing the newly obtained action, one can determine whether any non-dynamical perturbations remain by checking if all perturbations possess kinetic terms. In our newly obtained action, the perturbation $\phi_{h}$ leads to a constraint
$\phi_{h}$ $=$ $\Big(\kappa b_{t}\dot{H}k^{2}\dot{\varphi}_{b}+2\xi b_{t}H\dot{H}k^{2}\varphi_{b}-(\xi+2\sigma)(k^{2}-3a^{2}\dot{H})b_{t}H\dot{\psi}_{1}-\frac{3\kappa\overline{\rho}_{m,n}}{k^{2}}aH\dot{H}\dot{\phi}_{m}$ (96)
$-\big(\kappa\dot{H}k^{2}+\big((\xi-2\sigma)H^{2}-(\xi+2\sigma)\dot{H}\big)\big(k^{2}-3a^{2}\dot{H}\big)\big)b_{t}\psi_{1}+\Big(\big(\kappa k^{2}+3(\xi+2\sigma)a^{2}\dot{H}\big)a^{3}\dot{H}\overline{\rho}_{m,n}$
$-H^{2}(k^{2}-3a^{2}\dot{H})\big(4\sigma a^{3}\overline{\rho}_{m,n}+3(\xi+2\sigma)\overline{J}\overline{\rho}_{m,nn}\big)\Big)\frac{\kappa}{(\xi+2\sigma)a^{4}k^{2}}\phi_{m}\Big).$
@@
Substituting this constraint into the newly obtained action allows us to eliminate the variable $\phi_{h}$. Subsequently, we define two new variables to eliminate the coupling terms between the time derivatives of the remaining variables,
$\psi_{2}\equiv\varphi_{b}+\frac{(\xi+2\sigma)b_{t}^{2}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)H}\psi_{1}+\frac{3\kappa b_{t}a\dot{H}\overline{\rho}_{m,n}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)(k^{2}-3a^{2}\dot{H})Hk^{2}}\phi_{m},$ (97)
$\psi_{3}\equiv\psi_{1}-\frac{\kappa\overline{\rho}_{m,n}}{(\xi+2\sigma)b_{t}ak^{2}}\phi_{m}.$ (98)
@@
Using these two new variables, we can eliminate $\varphi_{b}$ and $\psi_{1}$ from the action. The resulting action can be written as
$S^{(2)}_{s}=\int dtd^{3}x\left[\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\dot{\psi}_{3}^{2}+\frac{a^{2}\overline{\rho}_{m,n}k^{2}}{2\overline{J}(k^{2}-3a^{2}\dot{H})}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}+\frac{aF_{1}(t,k)}{\kappa b_{t}^{2}\dot{H}k^{2}+2F_{1}(t,k)}(k\dot{\psi}_{2})^{2}+\cdots\right],$ (99)